# Deeper Insights into Deep Graph Convolutional Networks: Stability and Generalization

### v1

@@ Abstract
journal: IEEE Trans. Pattern Anal. Mach. Intell.
@@ 1 Introduction
Graph-structured structured data is pervasive across diverse domains, including knowledge graphs, traffic networks, and social networks to name a few [1, 2]. Several pioneering works [3, 4] introduced the initial concept of graph neural networks (GNNs), incorporating recurrent mechanisms and necessitating neural network parameters to define contraction mappings. Concurrently, Micheli [5] introduced the neural network for graphs, commonly referred to as NN4G, over a comparable timeframe. It is worth noting that the NN4G diverges from recurrent mechanisms and instead employs a feed-forward architecture, exhibiting similarities to contemporary GNNs. In recent years, (contemporary) GNNs have gained significant attention as an effective methodology for modeling graph data [6, 7, 8, 9, 10, 11]. To obtain a comprehensive understanding of GNNs and deep learning for graphs, we refer the readers to relevant survey papers for an extensive overview [12, 13, 14, 15].
@@
Among the various GNN variants, one of the most powerful and frequently used GNNs is graph convolutional networks (GCNs). A widely accepted perspective posits that GCNs can be regarded as an extension or generalization of traditional spatial filters, which are commonly employed in Euclidean data analysis, to the realm of non-Euclidean data. Due to its success on non-Euclidean data, GCN has attracted widespread attention on its theoretical exploration. Recent works on GCNs includes understanding over-smoothing [16, 17, 18, 19], interpretability and explainability[20, 21, 22, 23, 24], expressiveness [25, 26, 27], and generalization [28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. In this paper, we specifically address the generalization of GCNs to provide a bound on their generalization gap.
@@
Investigating the generalization of GCNs is essential in understanding its underlying working principles and capabilities from a theoretical perspective. However, the theoretical establishment in this area is still in its infancy. In recent work [36], Zhang et al. provided a novel technique based on algorithmic stability to investigate the generalization capability of single-layer GCNs in semi-supervised learning tasks. Their results indicate that the stability of a single-layer GCN trained with the stochastic gradient descent (SGD) algorithm is dependent on the largest absolute eigenvalue of graph filter operators. This finding highlights the crucial role of graph filters in determining the generalization capability of single-layer GCNs, providing guidance for designing effective graph filters for these networks. On the other hand, a number of prior studies have shown that deep GCNs possess greater expressive power than their single-layer counterparts. Consequently, it is essential to extend the generalization results of single-layer GCNs to their multi-layer counterparts. This will help us understand the effect of factors (e.g., graph filters, number of layers) on the generalization capability of deep GCNs.
In this paper, we study the generalization of deep GCNs. Our methods mainly follow the work proposed in [36] by estimating the uniform stability of the learning algorithm of deep GCNs in semi-supervised learning problems, but a more sophisticated analysis is required. The findings of our investigation reveal a strong association between the generalization gap of deep GCNs and the characteristics of the graph filter, particularly the number of layers employed. Specifically, we observe that if the maximum absolute eigenvalue (or the largest singular value) of graph filter operators remains invariant with respect to graph size, the generalization gap diminishes asymptotically at a rate of $O(1/\sqrt{m})$ as the training data size $m$ approaches infinity. This explains why normalized graph filters perform better than non-normalized ones in the deep GCN. Additionally, our results suggest that large number of layers can increase the generalization gap and subsequently degrade the performance of deep GCNs. This provides guidance for designing well-performing deep GCNs with a proper number of layers.
@@
    We prove the uniform stability of deep GCNs trained with SGD, which extends the findings of single-layer GCNs presented in [36].
@@
    An upper bound for the generalization gap of deep GCNs is provided with rigorous proofs. Our theoretical results shed light on the crucial components influencing the generalization ability of the deep GCN model.
@@
    Our empirical studies across three benchmark datasets for node classification verify convincingly our theoretical findings regarding the role of graph filters, the depth and width of deep GCN models.
The remainder of this paper is organized as follows. In Section 2, an overview of prior studies on the generalization of GCNs (or generic GNNs) is presented, along with a comparative analysis highlighting the similarities and distinctions between our work and previous research. Section 3 offers an exposition of the essential concepts. The primary findings of this paper are given in Section 4. Experimental studies designed to validate our theoretical findings are presented in Section 5. Section 6 concludes the paper with additional remarks. Detailed proofs of our theoretical results are included in the appendices.
@@ 2 Related Work
Theoretically, contemporary research on the generalization capability of GCNs predominantly employs methodologies such as Vapnik-Chervonenkis dimension (VC-dim) [30, 34], Rademacher complexity [31, 32, 33, 34, 35], and algorithmic stability [36, 37, 42, 43], as the mainstream categories revisited in this section. To provide a broader perspective, we also mention briefly other methodologies such as the classic PAC-Bayesian [38, 39], neural tangent kernels (NTK) [40, 41], algorithm alignment [44, 45], statistical physics and random matrix theory [46].
@@
VC-dim and Rademacher Complexity. In [30], Scarselli et al. examined the generalization capability of GNNs by providing upper bounds on the order of growth of the VC-dim of GNNs. While the VC-dim serves as a traditional concept for establishing learning bounds, its applicability does not account for the underlying graph structure. In [34], the authors also provided a generalization error bound for GNNs using VC-dim. However, the error bound based on VC-dim is trivial and fails to capture the beneficial impact of degree normalization. Esser et al. [34] explored the generalization upper bound using transductive Rademacher complexity (TRC), examining the impact of graph convolutions and network architectures on minimizing generalization errors and gaining insights into the conditions that enhance learning through the graph structure. Tang et al. [35] derived the upper bound for the generalization gap of popular GNNs by establishing high probability learning guarantees using transductive SGD. However, their upper bound depends on the dimensionality of parameters due to the inclusion of the parameter dimension in the TRC-based technique utilized for deriving the bounds.
Algorithmic Stability. In addition to VC-dim and Rademacher complexity, the uniform stability of learning algorithms plays a crucial role in the examination of generalization. Expanding on the previous discoveries made by Hardt et al. [47], Verma and Zhang [36] provided evidence that one-layer GCNs possess uniform stability characteristics and established an upper bound on generalization that scales in accordance with the largest absolute eigenvalue of the graph filter operator. In a continuation of the research presented in [36], Liu et al.[42] contributed to a comprehensive theoretical understanding of single-layer GCNs by analyzing the stability of their SGD proximal algorithm, incorporating $\ell_{p}$-regularization. However, it should be noted that these studies are limited to single-layer GCNs. Ng and Yip [37] focused their investigation on the stability and generalization properties of GCNs within eigen-domains. However, their formulation of the two-layer GCN relies on spectral graph convolution defined in [48], which necessitates the computationally expensive eigendecomposition of the graph Laplacian. As a result, their approach fails to offer meaningful theoretical insights for node classification in large-scale scenarios. In the context of this methodology, the studies most closely related to ours are [36] and [37]. However, unlike these works, our theoretical investigation specifically centers around deep GCNs without the constraint of assuming a spectral-based formulation.
Other Methodologies. The groundbreaking study conducted in [38] marks the initial endeavor to establish generalization bounds for GCNs and message passing neural networks, employing the PAC-Bayesian approach. Drawing upon an improved PAC-Bayesian analysis, the work presented in [39] establishes comprehensive generalization bounds for GNNs, highlighting a notable correlation with the graph diffusion matrix. Furthermore, the neural tangent kernel (NTK) introduced by [40] provides a promising avenue for investigating the generalization of infinitely wide multi-layer GNNs trained by gradient descent, as studied in [41]. These works however concentrate on the formulation of graph classification problems, instead of the node classification task (under a transductive setting) which poses a greater challenge. In addition, there exist related studies that employ a special theoretical framework distinct from ours, such as the analysis of the generalization capability of GNNs trained using topology-sampling techniques [49] or on large random graphs [50]. For a comprehensive review of the emerging theoretical perspectives on characterizing the capabilities of GNNs, we recommend interested readers to refer to [51].
@@ 3 Preliminaries and Notations
In this section, we provide a comprehensive description of the problem setup examined in this paper. Additionally, we present an extensive review of fundamental concepts related to uniform stability for training algorithms, which serve as the foundation for the subsequent analysis.
@@ 3.1 Deep Graph Convolutional Networks
$\mathbf{y}=\sigma\big{(}g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big{)},\$ (1)
@@
$\theta=\{\mathbf{W}^{(1)},\mathbf{W}^{(2)},\dots,\mathbf{W}^{(K)},\mathbf{w}\}.$
is a $K+1$ layers GCN with $K$ hidden layers and a final output layer, and in the case of $K=0$, it degenerates into the single-layer GCN studied in [36].
@@ 3.2 The SGD Algorithm
In this paper, $\mathcal{A}_{\mathcal{S}}$ is considered to be the algorithm given by the SGD algorithm. Following the approach employed in [36], our analysis focuses solely on the randomness inherent in $\mathcal{A}_{\mathcal{S}}$ arising from the SGD algorithm, while disregarding the stochasticity introduced by parameter initialization. The SGD algorithm for a deep GCN(1) aims to optimize its empirical error on a dataset $\mathcal{S}$ by updating parameters iteratively. For $t\in\mathbb{N}$ and considering the parameters $\theta_{t-1}$ obtained after $t-1$ iterations, the $t$-th iteration of SGD involves randomly drawing a sample $(\mathbf{x}_{t},y_{t})$ from the dataset $\mathcal{S}$. Subsequently, parameters $\theta$ are iteratively updated as follows:
@@ 3.3 Uniform Stability
For the sake of estimating the generalization gap $\epsilon_{gen}(\mathcal{A}_{\mathcal{S}})$ of $\mathcal{A}_{\mathcal{S}}$, we invoke the notion of uniform stability of $\mathcal{A}_{\mathcal{S}}$ as adopted in [52, 36].
@@
Definition 1 (Uniform Stability [36]).
@@
Lemma 1 (Stability Guarantees [36]).
@@ 4.1 Assumptions
With these assumptions, the derivative of $\sigma$, denoted by $\nabla\sigma$, is bounded, i.e., $|\nabla\sigma|\leq\alpha_{\sigma}$, and thus $\|\nabla\sigma(\mathbf{X})\|_{F}\leq\alpha_{\sigma}\|\mathbf{X}\|_{F}$ holds for any matrix $\mathbf{X}$. It can be easily verified that activation functions such as ELU and tanh satisfy the above assumptions.
Assumption 2. Let $\hat{y}$ and $y$ be the predicted and true labels, respectively. We denote the loss function $\ell:[y_{\min},y_{\max}]\times[y_{\min},y_{\max}]\rightarrow\mathbb{R}$ by $\ell(\hat{y},y)$. Similar to [37], we adopt the following assumptions for $\ell$.
@@
Table 1: Frequently used notations.
Notation Description
${\color[rgb]{0,0,0}g({\mathbf{L}})}$ the graph filter operator used in the considered deep GCNs
$C_{g}$ the 2-norm of $g(\mathbf{L})$, i.e., $C_{g}:=\|g(\mathbf{L})\|_{2}$
$C_{\mathbf{X}}$ the Frobenius norm of the input feature $\mathbf{\mathbf{X}}$, i.e., $C_{\mathbf{X}}:=\|\mathbf{X}\|_{F}$
$K$ the number of hidden layers of the considered deep GCNs
$\alpha_{\sigma}$, $\upsilon_{\sigma}$ parameters w.r.t the continuity of the activation function $\sigma(\cdot)$
$M$ the upper bound of the loss function $\ell(\cdot,\cdot)$
$m$ the number of samples in the trained dataset $\mathcal{S}$
$\eta$ the learning rate of $\mathcal{A}_{\mathcal{S}}$
$T$ the number of iterations for training $\mathcal{A}_{\mathcal{S}}$ using the SGD algorithm
$B$ the upper bound of the 2-norm of the parameters $\{\mathbf{W}^{(1)},\dots,\mathbf{W}^{(K)},\mathbf{w}\}$

### v2

@@ 1 Introduction
Graph-structured data is pervasive across diverse domains, including knowledge graphs, traffic networks, and social networks to name a few LIANG2022103714 , ma2021deep . Several pioneering works gori2005new , scarselli2008graph introduced the initial concept of graph neural networks (GNNs), incorporating recurrent mechanisms and necessitating neural network parameters to define contraction mappings. Concurrently, Micheli micheli2009neural introduced the neural network for graphs, commonly referred to as NN4G, over a comparable timeframe. It is worth noting that the NN4G diverges from recurrent mechanisms and instead employs a feed-forward architecture, exhibiting similarities to contemporary GNNs. In recent years, (contemporary) GNNs have gained significant attention as an effective methodology for modeling graph data yao2022multi , hamilton2020graph , GNNBook2022 , bianchi2021graph , jiang2023graph , zhang2024decouple . To obtain a comprehensive understanding of GNNs and deep learning for graphs, we refer the readers to relevant survey papers for an extensive overview bacciu2020gentle , Survey_SunMS , Survey_ZhangCQ , Survey_ZhuWW .
@@
Among the various GNN variants, one of the most powerful and frequently used GNNs is graph convolutional networks (GCNs). A widely accepted perspective posits that GCNs can be regarded as an extension or generalization of traditional spatial filters, which are commonly employed in Euclidean data analysis, to the realm of non-Euclidean data. Due to its success on non-Euclidean data, GCN has attracted widespread attention on its theoretical exploration. Recent works on GCNs includes understanding over-smoothing 2018Oversmoothing , 2020Oversmoothing , 2019GCN_expressive_power , 2020DropEdge , interpretability and explainability2020interpretability , 2021explainability , yuan2022explainability , schnake2021higher , bouritsas2022improving , expressiveness 2018how_powerful , 2019Approximation_GCNs , 2019understanding , and generalization 2019Du , 2020fastlearning_generalization , 2018VC-dim , 2020Rademacher , 2020RademacherOptimization , 2021Rademacher , 2021Rademacher_learning , 2023Generalization_GNNs , 2019Stable_GCN , 2023Stable_GCN , 2021PAC-generalization , 2023PAC-generalization , 2018_Neural-Tangent-Kernel , 2019_Neural-Tangent-Kernel . In this paper, we specifically address the generalization of GCNs to provide a bound on their generalization gap.
@@
Investigating the generalization of GCNs is essential in understanding its underlying working principles and capabilities from a theoretical perspective. However, the theoretical establishment in this area is still in its infancy. In recent work 2019Stable_GCN , Verma and Zhang provided a novel technique based on algorithmic stability to investigate the generalization capability of single-layer GCNs in semi-supervised learning tasks. Their results indicate that the stability of a single-layer GCN trained with the stochastic gradient descent (SGD) algorithm is dependent on the largest absolute eigenvalue of graph filter operators. This finding highlights the crucial role of graph filters in determining the generalization capability of single-layer GCNs, providing guidance for designing effective graph filters for these networks. On the other hand, a number of prior studies have shown that deep GCNs possess greater expressive power than their single-layer counterparts. Consequently, it is essential to extend the generalization results of single-layer GCNs to their multi-layer counterparts. This will help us understand the effect of factors (e.g., graph filters, number of layers) on the generalization capability of deep GCNs.
In this paper, we investigate the generalization properties of deep GCNs. Building on the stability framework of 2019Stable_GCN , we analyze the uniform stability of deep GCNs in semi-supervised learning, while developing a more refined theoretical treatment suited to deep architectures. Our analysis reveals a strong connection between the generalization gap of deep GCNs and the characteristics of the graph filter, particularly the number of layers. In particular, we show that when the maximum absolute eigenvalue (or the largest singular value) of the graph filter operator remains invariant with respect to graph size, the generalization gap diminishes asymptotically at a rate of $O(1/\sqrt{m})$ as the training sample size $m$ grows. This result explains why normalized graph filters generally outperform non-normalized ones in deep GCNs. Furthermore, our findings indicate that increasing depth can enlarge the generalization gap and consequently degrade performance, thereby offering theoretical guidance for selecting an appropriate number of layers when designing deep GCNs. We then empirically validate our theoretical results through experiments on three benchmark datasets: Cora, Citeseer, and Pubmed, demonstrating strong consistency between theory and practice. In addition, we further discuss how our theoretical framework extends to advanced architectures, including GCNII chen2020simple and Graph Transformer GT , thereby highlighting its broader applicability and its potential to inspire future theoretical studies on more complex GNN variants.
@@
    We establish the uniform stability of deep GCNs trained with SGD, thereby extending the earlier results on single-layer GCNs presented in 2019Stable_GCN .
@@
    We provide a rigorous upper bound for the generalization gap of deep GCNs and highlight the key factors that govern their generalization ability. Moreover, we further discuss how our theoretical framework extends naturally to advanced GNN architectures, including GCNII and Graph Transformer models.
@@
    We conduct empirical studies on three benchmark datasets for node classification, which strongly validate our theoretical findings regarding the influence of graph filters, as well as the depth and width of deep GCNs.
The remainder of this paper is organized as follows. In Section 2, an overview of prior studies on the generalization of GCNs (or generic GNNs) is presented, along with a comparative analysis highlighting the similarities and distinctions between our work and previous research. Section 3 offers an exposition of the essential concepts. The primary findings of this paper are given in Section 4. Experimental studies designed to validate our theoretical findings are presented in Section 5. In Section 6, we discuss how our findings extend to advanced GNN architectures, including GCNII and Graph Transformer models. Section 7 concludes the paper with additional remarks. The detailed proofs of our theoretical results are deferred to the Appendix section.
@@ 2 Related Work
Theoretical studies on the generalization capability of GCNs mainly employ three methodologies: Vapnik–Chervonenkis (VC) dimension 2018VC-dim , 2021Rademacher_learning , Rademacher complexity 2020Rademacher , 2020RademacherOptimization , 2021Rademacher , 2021Rademacher_learning , 2023Generalization_GNNs , and algorithmic stability 2019Stable_GCN , 2023Stable_GCN , 2023Stable_Lp , Litpami2023 . Other approaches include PAC-Bayesian theory 2021PAC-generalization , 2023PAC-generalization , neural tangent kernels (NTKs) 2018_Neural-Tangent-Kernel , 2019_Neural-Tangent-Kernel , algorithm alignment xu2019can , xu2020neural , and methods from statistical physics and random matrix theory shi2024homophily . For a broader perspective, we refer readers to the recent survey vasileiou2025survey , which provides a comprehensive overview of generalization theory for message-passing GNNs.
@@
VC-Dimension and Rademacher Complexity. Scarselli et al. 2018VC-dim study the generalization capability of GNNs by deriving upper bounds on the growth order of their VC-dimension. While VC-dimension is a classical tool for establishing learning bounds, it does not capture the structure of the underlying graph. Similarly, 2021Rademacher_learning provides VC-dimension–based error bounds for GNNs, but the results are trivial and fail to reflect the benefits of degree normalization. To address graph-specific effects, Esser et al. 2021Rademacher_learning analyze upper bounds using transductive Rademacher complexity (TRC), highlighting how graph convolutions and network architectures influence generalization. Tang et al. 2023Generalization_GNNs establish high-probability generalization bounds for popular GNNs via TRC-based analysis of transductive SGD. However, their bounds scale with the parameter dimension, limiting tightness for large models.
Algorithmic Stability. Beyond capacity-based measures, algorithmic stability serves as an important framework for understanding GNN generalization. Building on the work of Hardt et al. hardt2016train , Verma and Zhang 2019Stable_GCN show that one-layer GCNs exhibit uniform stability and provide generalization bounds that scale with the largest absolute eigenvalue of the graph filter operator. Extending this line, Liu et al. 2023Stable_Lp analyze the stability of single-layer GCNs trained with an SGD-proximal algorithm under $\ell_{p}$-regularization, yielding a more refined theoretical understanding. These studies, however, remain restricted to single-layer architectures. Cong et al. cong2021provable examine GNNs under uniform transductive stability, showing that deeper models improve stability and reduce generalization error, whereas our work adopts a different stability formulation. Ng and Yip 2023Stable_GCN investigate stability and generalization in two-layer GCNs under an eigen-domain formulation, relying on spectral graph convolution bruna2013spectral . Because this formulation requires computationally expensive eigendecomposition of the graph Laplacian, it does not scale to large node-classification tasks. Within this methodological line, the closest studies to ours are 2019Stable_GCN and 2023Stable_GCN , but our analysis focuses on deep GCNs without assuming a spectral-based formulation.
Other Methodologies. Alternative perspectives on GNN generalization also exist. The pioneering work of 2021PAC-generalization introduces PAC-Bayesian analysis for GCNs and message-passing neural networks, later extended in 2023PAC-generalization to provide tighter bounds linked to the graph diffusion matrix. The NTK framework introduced by 2018_Neural-Tangent-Kernel enables analysis of infinitely wide GNNs trained by gradient descent, with 2019_Neural-Tangent-Kernel extending this framework to multi-layer settings. However, NTK-based analyses typically focus on graph classification rather than the more challenging transductive node-classification setting. Additional work explores distinct theoretical frameworks, including topology-sampling techniques li2022generalization , analysis on large random graphs keriven2020convergence , and NTK-based loss landscape analysis of wide GCNs zhou2025tighter . For further perspectives, we refer readers to the survey jegelka2022theory , which synthesizes emerging theoretical approaches to characterizing GNN capabilities.
@@ 3 Preliminaries and Notations
In this section, we describe the problem setup considered in this paper and review fundamental concepts of uniform stability for training algorithms, which form the basis of our subsequent analysis. For clarity, we first summarize the main symbols used in this paper in the table below.
Table 1: Frequently used notations.
Notation Description
$\color[rgb]{0,0,0}g({\mathbf{L}})$} graph filter operator used in the considered deep GCNs
$C_{g}$ the 2-norm of $g(\mathbf{L})$, i.e., $C_{g}:=\|g(\mathbf{L})\|_{2}$
$C_{\mathbf{X}}$ Frobenius norm of the input feature $\mathbf{\mathbf{X}}$, i.e., $C_{\mathbf{X}}:=\|\mathbf{X}\|_{F}$
$K$ number of hidden layers of the considered deep GCNs
$\alpha_{\sigma}$, $\upsilon_{\sigma}$ parameters w.r.t the continuity of activation function $\sigma(\cdot)$
$\nabla\sigma$ the derivative of activation function $\sigma(\cdot)$
$M$ the upper bound of loss function $\ell(\cdot,\cdot)$
$m$ the number of samples in the trained dataset $\mathcal{S}$
$\eta$ the learning rate of $\mathcal{A}_{\mathcal{S}}$
$T$ number of iterations for training $\mathcal{A}_{\mathcal{S}}$ using SGD
$\mu_{m}$ the uniform stability of a learning algorithm $\mathcal{A}_{\mathcal{S}}$
$\boldsymbol{\delta}_{\mathbf{x}}$ the indicator vector with respect to node $\mathbf{x}$
$\boldsymbol{\delta}_{i}$ the indicator vector with respect to index $i$
$\mathbf{X}^{(k)}$ the output feature matrix of the $k$-th layer
$\triangle\mathbf{X}^{(k)}$ the variation of $\mathbf{X}^{(k)}$ in two GCNs
$\mathbf{W}^{(k)}$ the parameter matrix specific to the $k$-th layer
$B$ upper bound for 2-norm of $\{\mathbf{W}^{(1)},\dots,\mathbf{W}^{(K)},\mathbf{w}\}$
$\triangle\mathbf{W}^{(k)}$ the variation of $\mathbf{W}^{(k)}$ in two GCNs
$\triangle\theta$ $\triangle\theta=\{\triangle\mathbf{W}^{(1)},\dots,\triangle\mathbf{W}^{(K)},\triangle\mathbf{w}\}$
$\mathbf{W}^{(k)}_{t}$ the learnt $\mathbf{W}^{(k)}$ trained after $t$ iterations
$\triangle\mathbf{W}^{(k)}_{t}$ the variation of $\mathbf{W}^{(k)}_{t}$ of two GCNs trained after $t$ iterations
$\triangle\theta_{t}$ $\triangle\theta_{t}=\{\triangle\mathbf{W}_{t}^{(1)},\dots,\triangle\mathbf{W}_{t}^{(K)},\triangle\mathbf{w}_{t}\}$
@@ 3.1 Deep Graph Convolutional Networks
$\mathbf{y}=\sigma\big(g(\mathbf{L})\mathbf{X}^{(K)}\mathbf{w}\big),\\$ (1)
@@
$\theta=\{\mathbf{W}^{(1)},\mathbf{W}^{(2)},\dots,\mathbf{W}^{(K)},\mathbf{w}\}$
is a $K+1$ layers GCN with $K$ hidden layers and a final output layer, and in the case of $K=0$, it degenerates into the single-layer GCN studied in 2019Stable_GCN .
@@ 3.2 The SGD Algorithm
In this paper, $\mathcal{A}_{\mathcal{S}}$ is considered to be the algorithm given by the SGD algorithm. Following the approach employed in 2019Stable_GCN , our analysis focuses solely on the randomness inherent in $\mathcal{A}_{\mathcal{S}}$ arising from the SGD algorithm, while disregarding the stochasticity introduced by parameter initialization. The SGD algorithm for a deep GCN (1) aims to optimize its empirical error on a dataset $\mathcal{S}$ by updating parameters iteratively. For $t\in\mathbb{N}$ and considering the parameters $\theta_{t-1}$ obtained after t1t-1 iterations, the $t$-th iteration of SGD involves randomly drawing a sample $(\mathbf{x}_{t},y_{t})$ from the dataset $\mathcal{S}$. Subsequently, parameters $\theta$ are iteratively updated as follows:
@@ 3.3 Uniform Stability
For the sake of estimating the generalization gap $\epsilon_{gen}(\mathcal{A}_{\mathcal{S}})$ of $\mathcal{A}_{\mathcal{S}}$, we invoke the notion of uniform stability of $\mathcal{A}_{\mathcal{S}}$ as adopted in 2005stability_random , 2019Stable_GCN .
@@
Definition 1 (Uniform Stability 2019Stable_GCN ).
@@
Lemma 1 (Stability Guarantees 2019Stable_GCN ).
@@ 4.1 Assumptions
With these assumptions, the derivative of $\sigma$, denoted by $\nabla\sigma$, is bounded, i.e., $|\nabla\sigma(\cdot)|\leq\alpha_{\sigma}$, and $\|\sigma(\mathbf{X})\|_{F}\leq\alpha_{\sigma}\|\mathbf{X}\|_{F}$ holds for any matrix $\mathbf{X}$. It can be easily verified that activation functions such as ELU and tanh satisfy the above assumptions.
Assumption 2. Let $\hat{y}$ and $y$ be the predicted and true labels, respectively. We denote the loss function $\ell:[y_{\min},y_{\max}]\times[y_{\min},y_{\max}]\rightarrow\mathbb{R}$ by $\ell(\hat{y},y)$. Similar to 2023Stable_GCN , we adopt the following assumptions for $\ell$.